OpenStudy (anonymous):

(4x)/(x+2)+(3)/(x-1)=(15)/(x^2+3x-4) help please?

OpenStudy (anonymous):

\[ \frac{ 4x }{ x+2}+\frac{ 3 }{ x-1 }=\frac{ 15 }{ x ^{2}+3x-4}\]

OpenStudy (anonymous):

add the two fractions on the LEFT then cross multiply...

OpenStudy (anonymous):

wait....but there's a equal sign between them though?

OpenStudy (anonymous):

yeah... when you have equal proportions you can do that... for example: \(\huge \frac{a}{b}=\frac{c}{d} \) cross multiplying yields: \(\huge a\cdot d = b \cdot c \)

OpenStudy (anonymous):

oh okay thankyou(: i got it<3

OpenStudy (anonymous):

yw...:)